# Dynamic Sparse Training versus Dense Training: The Unexpected Winner in Image Corruption Robustness

### v1

@@ 6 Further Analysis and Discussion
6.1 DSCR Hypothesis in Recent DST Algorithms
Table 2: Robustness accuracy (%) on Fashion MNIST-C.
Dense Dynamic Sparse Training
Training Sparsity SET RigL CHTs+CSTI CHTs+BSW
95% 25.55 $\pm$ 0.32 25.03 $\pm$ 1.11 27.38 $\pm$ 0.25 27.13 $\pm$ 0.33
In this paper, we primarily investigate the robustness of fundamental DST models initialized from random sparse networks with the basic removal and regrowth criteria, to validate our DSCR hypothesis in commonly used settings. With the emergence of other recent and promising DST methods, we also include an initial exploration to further validate our hypothesis on these methods, such as Cannistraci-Hebb soft training (CHTs) Zhang et al. (2024b; a), which employ more complex growth criteria and topology initialization. We compare dense training with CHTs using Correlated Sparse Topological Initialization (CHTs+CSTI) and bipartite small-world networks (CHTs+BSW) on robustness accuracy, using the Fashion-MNIST corrupted dataset with an MLP model, as shown in Table 2. The results indicate that the recent DST method (i.e., CHTs) achieves even more inspiring robustness performance, further validating our DSCR hypothesis on the robustness of DST.
6.2 Discussion and Limitation
Our study questions the conventional approach of using ”de facto” dense training for model robustness in favor of sparse training. Due to the page constraint, we also analyze the complementarity between DST and data augmentation (e.g., Mixup Zhang et al. (2018)) in Appendix A.6, and further explore DST’s robustness with Out-Of-Domain (OOD) data and Grad-CAM Selvaraju et al. (2017) in Appendix A.7 and A.8, respectively. In the future, it is worth moving beyond the basic fact revealed in this paper, i.e., vanilla DST outperforms vanilla Dense Training for model robustness, and to devise new DST methods for model robustness especially in the high sparsity regime, and to replace Dense Training with DST in the current state-of-the-art techniques for model robustness. Additionally, while we empirically analyze robustness behaviors between dense and sparse training and provide insights from both spatial and frequency domains, further work on the theoretical understanding of this phenomenon, along with a fine-grained exploration of different components of DST (e.g., topology initialization, regrowth methods and regrowth weight initialization), could provide deeper insights for the development of more robust and efficient networks. This would be a promising direction for future research.
@@ 7 Conclusion
This paper questioned the typical approach of using ”de facto” Dense Training to maximize model robustness in deep learning, and revealed an unexpected finding. That is: ”Dynamic Sparse Training at low sparsity levels can train more robust models against image corruption than Dense Training.”. This finding has been confined formally in the DSCR Hypothesis and rigorously validated on nine scenarios (as summarized in Table 3 where a striking observation can be made - all DST algorithms studied outperform Dense Training in all scenarios). To understand what factors contribute to this compelling and intriguing DST behaviour, to the end, we present a deeper analyze. We show that the dynamic sparsity process performs a form of implicit regularization which makes the model to automatically learn to focus on the more informative features (e.g., rich in low-frequency), while paying less attention to the less informative ones (e.g., rich in high-frequency).
@@ A.2 Related benchmarks
  • Tiny-ImageNet-C Hendrycks & Dietterich (2019) consists of 15 types of common image corruptions, and is synthetically generated from Tiny-ImageNet dataset as well. The types of corruption are Gaussian noise, impulse noise, shot noise, defocus blur, glass blur, motion blur, zoom blur, brightness, fog, frost, snow, contrast, pixelate, elastic transform, and JPEG compression, with 5 severity levels.
  • ImageNet-C Hendrycks & Dietterich (2019) consists of 15 types of common image corruptions, and is synthetically generated from Tiny-ImageNet dataset as well. The types of corruption are Gaussian noise, impulse noise, shot noise, defocus blur, glass blur, motion blur, zoom blur, brightness, fog, frost, snow, contrast, pixelate, elastic transform, and JPEG compression, with 5 severity levels.
@@ A.3.1 Implementation Details
For CIFAR10, CIFAR100, and TinyImageNet, the codes were implemented in Python using PyTorch and executed on a single NVIDIA A100 GPU. Each was run three times using three fixed random seeds. The ImageNet experiments were run using distributed computing on four NVIDIA A100 GPUs. A batch size of 64 was utilized for each task, resulting in a combined total batch size of 256.
Table 4: The experiment hyperparameters setting. The hyperparameters include Learning Rate (LR), Batch Size (BS), LR Schedule (LRS), Optimizer (Opt), Weight Decay (WD), Sparsity Distribution (Sparsity Dist), Topology Update Interval ($\Delta T$), Pruning Rate Schedule (Sched), Initial Pruning Rate (P).
@@ A.3.2 The Computational and Memory Consumption
Table 5: Training FLOPs (TrFLOPs, ×$\times 10^{16}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for VGG16 on CIFAR10 at various sparsity ratios ($S$) using different sparsified algorithms.
TrFLOPs (×$\times 10^{16}$) TeFLOPs (×$\times 10^{8}$) Param (×$\times 10^{6}$)
$density$ 0.3 0.4 0.5 0.6 0.7 1.0 0.3 0.4 0.5 0.6 0.7 1.0 0.3 0.4 0.5 0.6 0.7 1.0
@@
AC/DC 0.89 1.03 1.11 1.19 1.27 - 1.92 2.55 3.17 3.80 4.42 - 4.95 6.43 7.89 9.37 10.81 -
CIFAR-100: For the CIFAR100 dataset, we train ReNet-34 for 160 epochs. The soft memory bound for MEST is set to 10% of the target density. For example, if the sparse network’s density is 0.3, then the soft memory bound is 0.03, resulting in an initial model density of 0.33. This soft memory bound is decayed to 0 using a polynomial strategy. The initial density of GraNet is set at 0.8. Every 500 training steps, it prunes weights to achieve a density described by the formula: $d_{i}-(d_{i}-d_{t})\left(1-\frac{t-t_{0}}{n\Delta t}\right)^{3}$, where $d_{i}$ is the initial density, $d_{t}$ is the target density, and $n\Delta t$ is set to half of the total training time.
Table 6: Training FLOPs (TrFLOPs, ×$\times 10^{16}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for ResNet34 on CIFAR100 at various sparsity ratios ($S$) using different sparsified algorithms.
@@
Table 7: Training FLOPs (TrFLOPs, ×$\times 10^{15}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for EfficientNet-B0 on TinyImageNet at various sparsity ratios ($S$) using different sparsified algorithms.
@@
Table 8: Training FLOPs (TrFLOPs, ×$\times 10^{18}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for ResNet50 on ImageNet at various sparsity ratios ($S$) using different sparsification algorithms.
@@
Figure 9: The comparison of training FLOPs, parameter count (represented by the area of each circle), and robustness accuracy of the dense and dynamic sparse EfficientNet-B0 on TinyImageNet. Each point represents the result for a specific sparsity ratio.
@@ A.4 The robustness accuracy for different corruption types and serverities
Figure 10: Compared with the the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTr, (c) GraNetr under different severities of CIFAR10-C.
Figure 11: Compared with the the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTg, (c) GraNetg under different severities of CIFAR10-C.
Figure 12: Compared with the the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTr, (c) GraNetr under different severities of CIFAR100-C.
Figure 13: Compared with the the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTg, (c) GraNetg under different severities of CIFAR100-C.
Figure 14: Compared with the the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTr, (c) GraNetr under different severities of TinyImageNet-C.
Figure 15: Compared with the the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTg, (c) GraNetg under different severities of TinyImageNet-C.
@@ A.5 Analysis Through the Lens of Filter
Figure 16: Visualizing non-zero weight count (1st row) and the sum of weight magnitudes (2nd row) in filters of VGG16 (#layer feature.20, #layer feature.30 and #layer feature.40 with kernel size $3\times 3$) after training on CIFAR100 using different sparsification algorithms: (b) SET, (c) RigL, (d) MESTr and (e) GraNetr, compared with (a) dense counterpart. Each point represents the corresponding value of a filter.
Figure 17: Visualizing non-zero weight count (1st row) and the sum of weight magnitudes (2nd row) in filters of ResNet34 (#layer2.0.conv2, #layer2.1.conv1 and #layerlayer3.3.conv2 with kernel size $3\times 3$) after training on CIFAR100 using different sparsification algorithms: (b) SET, (c) RigL, (d) MESTr and (e) GraNetr, compared with (a) dense counterpart. Each point represents the corresponding value of a filter.
@@
A.6 How data augmentation interacts with DST?
Table 10: Robustness accuracy (%) on CIFAR100 dataset with and without Mixup.
w/ mixup Dense training SET
$\times$ 51.68 52.60
$\checkmark$ 54.75 54.93
A.7 The performance of DST on out-of-domain (OOD) test data
@@
A.8 Grad-CAM
Grad-CAM (Gradient-weighted Class Activation Mapping) Selvaraju et al. (2017) generates a heatmap highlighting important regions by weighting the feature maps based on the gradient values. It provides practitioners with a visual indication of the areas on which the network focuses when making predictions. We visualize and compare the Grad-CAM outputs of both dense models and dynamic sparse models, as shown in Figure 19. We present different cases, including clean background ((a) and (b)), and complex background (c). Interestingly, we observe that the highlighting region of dynamic sparse models is smaller than that of the dense model. In Figure 19 (a), the dynamic sparse models focus on the face of the cat for prediction, while the dense models pay attention to both the face and body of the cat. In Figure 19 (b), the dynamic sparse models specifically emphasize the wings of the eagle. In contrast, the dense models primarily concentrate on a single wing while also capturing irrelevant contexts, leading to larger attention regions. In the case of a more complex background (In Figure 19 (c)), the dynamic sparse models tend to allocate some attention to the surrounding elements. However, the attention regions are still relatively smaller and primarily focused on the face of the panda, especially in the sparse model sparsified by SET.

### v2

@@ 6 Further Analysis and Discussion
Our study questions the conventional approach of using “de facto” dense training for model robustness in favor of sparse training. Due to the page constraint, we also analyze the complementarity between DST and data augmentation (e.g., Mixup Zhang et al. (2018)) in Appendix A.7, and further explore DST’s robustness with Out-Of-Domain (OOD) data and Grad-CAM Selvaraju et al. (2017) in Appendix A.8 and A.10, respectively. Furthermore, we extend and validate our DSCR hypothesis with more recent DST methods, including Cannistraci-Hebb soft training (CHTs) Zhang et al. (2024b; a), which employs more complex growth criteria and topology initialization, as detailed in Appendix A.6. In the future, it is worth moving beyond the basic fact revealed in this paper, i.e., vanilla DST outperforms vanilla Dense Training for model robustness, and to devise new DST methods for model robustness especially in the high sparsity regime, and to replace Dense Training with DST in the current state-of-the-art techniques for model robustness. Additionally, while we empirically analyze robustness behaviors between dense and sparse training and provide insights from both spatial and frequency domains, further work on the theoretical understanding of this phenomenon, along with a fine-grained exploration of different components of DST (e.g., topology initialization, regrowth methods and regrowth weight initialization), could provide deeper insights for the development of more robust and efficient networks. This would be a promising direction for future research.
@@ 7 Conclusion
This paper questioned the typical approach of using “de facto” Dense Training to maximize model robustness in deep learning, and revealed an unexpected finding. That is: “Dynamic Sparse Training at low sparsity levels can train more robust models against image corruption than Dense Training.”. This finding has been confined formally in the DSCR Hypothesis and rigorously validated on nine scenarios (as summarized in Table 2 where a striking observation can be made - all DST algorithms studied outperform Dense Training in all scenarios). To understand what factors contribute to this compelling and intriguing DST behaviour, to the end, we present a deeper analyze. We show that the dynamic sparsity process performs a form of implicit regularization which makes the model to automatically learn to focus on the more informative features (e.g., rich in low-frequency), while paying less attention to the less informative ones (e.g., rich in high-frequency).
Acknowledgements
Qiao Xiao is supported by the research program ‘MegaMind - Measuring, Gathering, Mining and Integrating Data for Self-management in the Edge of the Electricity System’, (partly) financed by the Dutch Research Council (NWO) through the Perspectief program under number P19-25. Elena Mocanu is partly supported by the Modular Integrated Sustainable Datacenter (MISD) project funded by the Dutch Ministry of Economic Affairs and Climate under the European Important Projects of Common European Interest - Cloud Infrastructure and Services (IPCEI-CIS) program for 2024-2029. This research used the Dutch national e-infrastructure with the support of the SURF Cooperative, using grant no. EINF-7499.
@@ A.2 Related benchmarks
  • Tiny-ImageNet-C Hendrycks & Dietterich (2019) consists of 15 types of common image corruptions, and is synthetically generated from Tiny-ImageNet dataset. The types of corruption are Gaussian noise, impulse noise, shot noise, defocus blur, glass blur, motion blur, zoom blur, brightness, fog, frost, snow, contrast, pixelate, elastic transform, and JPEG compression, with 5 severity levels.
  • ImageNet-C Hendrycks & Dietterich (2019) consists of 15 types of common image corruptions, and is synthetically generated from ImageNet dataset. The types of corruption are Gaussian noise, impulse noise, shot noise, defocus blur, glass blur, motion blur, zoom blur, brightness, fog, frost, snow, contrast, pixelate, elastic transform, and JPEG compression, with 5 severity levels.
@@ A.3.1 Implementation Details
For CIFAR10, CIFAR100, and TinyImageNet, the codes were implemented in Python using PyTorch and executed on a single NVIDIA A100 GPU. Each was run three times using three fixed random seeds. The ImageNet experiments were run using distributed computing on four NVIDIA A100 GPUs. A batch size of 64 was utilized for each task, resulting in a combined total batch size of 256. For the UCF101 dataset, we followed the settings described in Schiappa et al. (2023), utilizing the codebase  [^5] https://github.com/Maddy12/ActionRecognitionRobustnessEval. For the FashionMNIST dataset, we followed the training settings detailed in Zhang et al. (2024b; a), using the codebase  [^6] https://github.com/biomedical-cybernetics/Cannistraci-Hebb-training.
Table 3: The experimental hyperparameters settings. The hyperparameters include Learning Rate (LR), Batch Size (BS), LR Schedule (LRS), Optimizer (Opt), Weight Decay (WD), Sparsity Distribution (Sparsity Dist), Topology Update Interval ($\Delta T$), Pruning Rate Decay Schedule (Sched), Initial Pruning Rate (P).
@@ A.3.2 The Computational and Memory Consumption
Table 4: Training FLOPs (TrFLOPs, ×$\times 10^{16}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for VGG16 on CIFAR10 at various sparsity ratios ($S$) using different DST algorithms.
TrFLOPs (×$\times 10^{16}$) InFLOPs (×$\times 10^{8}$) Param (×$\times 10^{6}$)
$S$ 0.7 0.6 0.5 0.4 0.3 0.0 0.7 0.6 0.5 0.4 0.3 0.0 0.7 0.6 0.5 0.4 0.3 0.0
@@
CIFAR-100: For the CIFAR100 dataset, we train ResNet34 for 160 epochs. The soft memory bound for MEST is set to 10% of the target density. For example, if the sparse network’s density is 0.3, then the soft memory bound is 0.03, resulting in an initial model density of 0.33. This soft memory bound is decayed to 0 using a polynomial strategy. The initial density of GraNet is set at 0.8. Every 500 training steps, it prunes weights to achieve a density described by the formula: $d_{i}-(d_{i}-d_{t})\left(1-\frac{t-t_{0}}{n\Delta t}\right)^{3}$, where $d_{i}$ is the initial density, $d_{t}$ is the target density, and $n\Delta t$ is set to half of the total training time.
Table 5: Training FLOPs (TrFLOPs, ×$\times 10^{16}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for ResNet34 on CIFAR100 at various sparsity ratios ($S$) using different DST algorithms.
@@
Table 6: Training FLOPs (TrFLOPs, ×$\times 10^{15}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for EfficientNet-B0 on TinyImageNet at various sparsity ratios ($S$) using different DST algorithms.
@@
Table 7: Training FLOPs (TrFLOPs, ×$\times 10^{18}$), Inference FLOPs (InFLOPs, ×$\times 10^{9}$), and the number of model parameters (Param, ×$\times 10^{6}$) for ResNet50 on ImageNet at sparsity ratios ($S=0.1$) using different DST algorithms.
@@
Figure 9: The comparison of training FLOPs, parameter count (represented by the area of each circle), and robustness accuracy of the dense and sparse EfficientNet-B0 on TinyImageNet. Each point represents the result for a specific sparsity ratio.
@@
It is worth noting that the computational costs in this work are theoretical and not fully realized gains on current hardware infrastructures. This is a common limitation in the literature on sparse training methods, as most current hardware and software systems are optimized for dense matrix operations rather than sparse ones. However, recent advancements in model sparsity are increasingly aligning with hardware and software developments to fully leverage the benefits of sparsity. For example, NVIDIA’s A100 GPU supports 2:4 sparsity Zhou et al. (2021), and other innovations are making strides toward efficient sparse implementations Chen et al. (2019); Ashby et al. (2019). Simultaneously, software libraries such as Liu et al. (2021b); Mocanu et al. (2018) are emerging to enable truly sparse network implementations. These advancements are paving the way for future deep neural networks to achieve greater efficiency in terms of computation, memory, and energy use.
@@ A.4 The robustness accuracy for different corruption types and serverities
Figure 10: Compared with the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) SET, (b) MESTr, (c) GraNetr under different severities of CIFAR10-C.
Figure 11: Compared with the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTg, (c) GraNetg under different severities of CIFAR10-C.
Figure 12: Compared with the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) SET, (b) MESTr, (c) GraNetr under different severities of CIFAR100-C.
Figure 13: Compared with the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTg, (c) GraNetg under different severities of CIFAR100-C.
Figure 14: Compared with the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) SET, (b) MESTr, (c) GraNetr under different severities of TinyImageNet-C.
Figure 15: Compared with the dense model, the relative robustness gain of different dynamic sparse models with sparsity $0.5$ trained by (a) RigL, (b) MESTg, (c) GraNetg under different severities of TinyImageNet-C.
@@ A.5 Analysis Through the Lens of Filter
Figure 16: Visualizing non-zero weight count (1st row) and the sum of weight magnitudes (2nd row) in filters of VGG16 (#layer feature.20, #layer feature.30 and #layer feature.40 with kernel size $3\times 3$) after training on CIFAR10 using different DST algorithms: (b) SET, (c) RigL, (d) MESTr and (e) GraNetr, compared with (a) dense counterpart. Each point represents the corresponding value of a filter.
Figure 17: Visualizing non-zero weight count (1st row) and the sum of weight magnitudes (2nd row) in filters of ResNet34 (#layer2.0.conv2, #layer2.1.conv1 and #layerlayer3.3.conv2 with kernel size $3\times 3$) after training on CIFAR100 using different DST algorithms: (b) SET, (c) RigL, (d) MESTr and (e) GraNetr, compared with (a) dense counterpart. Each point represents the corresponding value of a filter.
@@
Table 8: Robustness accuracy (%) of MLP on Fashion MNIST-C.
Dense Dynamic Sparse Training
Training Sparsity SET RigL CHTs+CSTI CHTs+BSW
95% 25.55 $\pm$ 0.32 25.03 $\pm$ 1.11 27.38 $\pm$ 0.25 27.13 $\pm$ 0.33
A.6 DSCR Hypothesis in Recent DST Algorithms
In this paper, we primarily investigate the robustness of fundamental DST models initialized from random sparse networks with the basic removal and regrowth criteria, to validate our DSCR hypothesis in commonly used settings. With the emergence of other recent and promising DST methods, we also include an initial exploration to further validate our hypothesis on these methods, such as Cannistraci-Hebb soft training (CHTs) Zhang et al. (2024b; a), which employ more complex growth criteria and topology initialization. We compare dense training with CHTs using Correlated Sparse Topological Initialization (CHTs+CSTI) and bipartite small-world networks (CHTs+BSW) on robustness accuracy, using the Fashion-MNIST corrupted dataset with an MLP model, as shown in Table 8. More training details are provided in Appendix A.3.1. The results indicate that the recent DST method (i.e., CHTs) achieves even more inspiring robustness performance, further validating our DSCR hypothesis on the robustness of DST.
A.7 How data augmentation interacts with DST?
Table 9: Robustness accuracy (%) on CIFAR100 with and without Mixup.
w/ mixup Dense training SET
$\times$ 51.68 52.60
$\checkmark$ 54.75 54.93
A.8 The performance of DST on out-of-domain (OOD) test data
@@
A.9 The performance of DST on the segmentation task
We compared the performance of a dense trained fully convolutional network (FCN) with a VGG16 backbone on the Pascal VOC 2012 Everingham et al. clean dataset, then test on corrupted Pascal VOC 2012 (generated following the ImageCorruption library  [^8] https://github.com/bethgelab/imagecorruptions, which includes 19 types of corruption at 5 severity levels). The results are presented in Table 12.
Table 12: Clean metric (%) on Pascal VOC 2012 and Robustness metric (%) on corrupted Pascal VOC 2012.
Metric Dense training SET (s=0.2) SET (s=0.1)
clean pixAcc ($\uparrow$) 85.65 85.63 85.66
clean mIoU ($\uparrow$) 46.34 46.69 47.10
robustness pixAcc ($\uparrow$) 76.80 77.30 77.39
robustness mIoU ($\uparrow$) 22.60 22.73 23.09
Unlike classification tasks on CIFAR or ImageNet, where the sparse topology is randomly initialized, we first prune the backbone to obtain the sparse model, then perform sparse fine-tuning on the training set of the Pascal VOC 2012 clean dataset, and finally test it on both the clean and corrupted versions of the Pascal VOC 2012 test set, as the VGG16 backbone for the segmentor is typically pretrained on ImageNet. The results show that DST models are not only robust in image and video classification but also in image segmentation tasks.
A.10 Grad-CAM
Grad-CAM (Gradient-weighted Class Activation Mapping) Selvaraju et al. (2017) generates a heatmap highlighting important regions by weighting the feature maps based on the gradient values. It provides practitioners with a visual indication of the areas on which the network focuses when making predictions. We visualize and compare the Grad-CAM outputs of both dense models and dynamic sparse models, as shown in Figure 19. We present different cases, including clean background ((a) and (b)), and complex background (c). Interestingly, we observe that the highlighting region of dynamic sparse models is smaller than that of the dense model. In Figure 19 (a), the dynamic sparse models focus on the face of the cat for prediction, while the dense models pay attention to both the face and body of the cat. In Figure 19 (b), the dynamic sparse models specifically emphasize the wings of the eagle. In contrast, the dense models primarily concentrate on a single wing while also capturing irrelevant contexts, leading to larger attention regions. In the case of a more complex background (In Figure 19 (c)), the dynamic sparse models tend to allocate some attention to the surrounding elements. However, the attention regions are still relatively smaller and primarily focused on the face of the bear, especially in the sparse model sparsified by SET.